OpenStudy (joshoyen):

Factor our the common factor first. A. 2n^2 - 16 = 0 B. 7n^2 - 21 = 0

pooja195 (pooja195):

A What is a common factor of 2 and 16? (greatest)

OpenStudy (joshoyen):

2

pooja195 (pooja195):

Ok so now when we divide both if the numbers by 2 what are we left with?

OpenStudy (joshoyen):

n^2 - 8 = 0

OpenStudy (joshoyen):

hm, na thats what im stuck on

pooja195 (pooja195):

Wait i messed up sorry >.<

OpenStudy (joshoyen):

oh lol , no problem

pooja195 (pooja195):

You have the common factor correct But before the we need to add 16 to both sides

OpenStudy (joshoyen):

oh, okay

OpenStudy (joshoyen):

2n^2 = 16

pooja195 (pooja195):

\[\huge~\rm~2n^2 - 16+16 = 0+16\] \[\huge~\rm~2n^2 =16\]

OpenStudy (joshoyen):

yep

pooja195 (pooja195):

Now we divide both sides by 2

OpenStudy (joshoyen):

n^2 = 8

pooja195 (pooja195):

\[\huge~\rm~\sqrt{n^2} = \sqrt{8}\]

OpenStudy (joshoyen):

aight now i have to find the square root of 8

pooja195 (pooja195):

There is no square root of 8...

pooja195 (pooja195):

What exactly is your question??

OpenStudy (danjs):

Refer to the last prob i helped with, same to solve after factoring

OpenStudy (joshoyen):

ohhkay, nvm thanks

OpenStudy (danjs):

remember this thing \[\sqrt{x^2}=\left| x \right|\]

OpenStudy (joshoyen):

yeah

OpenStudy (danjs):

and the other side is root of 8, and so x can be + or - root of 8

OpenStudy (joshoyen):

alright thanks

OpenStudy (danjs):

only difference is they want you to factor first, then divide and solve.. instead of just moving the 16 over and dividing by 2, then root everything,

OpenStudy (danjs):

same thing really, you are multiplying the left by 2/2 or 1, which does not change it

OpenStudy (danjs):

(2/2) * (2n^2 - 16) = 0

OpenStudy (danjs):

you left the top 2 out front, and divided each term by 2, (multiply by 1/2) to get that factored form ..

OpenStudy (joshoyen):

ohh, okay thanks man

OpenStudy (danjs):

welcome, i hope ya see er clearly

OpenStudy (joshoyen):

yep i got it

OpenStudy (danjs):

the factored form is then 2*(n^2 - 8) = 0 that is all they want you to show, then solve normally, divide by 2, add 8, root

OpenStudy (joshoyen):

yep , gotcha thanks a lot man